OpenStudy (anonymous):

Solve the following system of linear equations by substitution: 6x+6y=12 3x+3y=6

OpenStudy (agreene):

All points along: y=2-x are solutions. 6x+6y=12 ==> y=2-x 3x+3x=6 ==>y=2-x

OpenStudy (agreene):

I did what I wrote, I solved both in terms of y and found they are the same thing... you could tell this was going to be the solution because both sets of equations had equal multiples of the intercept as their binary variable constants

OpenStudy (anonymous):

@merlynn It's a pleasure to help person who will say "Thank you" and give a medal for the helpful answer.

OpenStudy (anonymous):

Thank you.